Find the sum to n terms of the series 1•2•3+2•4•6+3•6•9+.........

Grade:10

1 Answers

Ojas Raundale
7 years ago
1.2.3 + 2.4.6 + 3.6.9 +.....+ n.2n.3n= 1.2.3[ 1^3 + 2^3 +3^3 +4^3+ .... + n^3]=1.2.3[ (n)(n+1)/2]^2=3/2[(n)(n+1)]^2